OpenStudy (itrymath):

A school system is reducing the amount of dumpster loads of trash removed each week. In week 5, there were 45 dumpster loads of waste removed. In week 10, there were 30 dumpster loads removed. Assume that the reduction in the amount of waste each week is linear. Write an equation in function form to show the amount of trash removed each week. (2 points) f(x) = 3x + 45 f(x) = - 3x + 60 f(x) = - 3x + 45 f(x) = 3x + 60

OpenStudy (itrymath):

Hey before we get started but i probably know answer.... do you know anything about forex, or profits, trade etc...

Directrix (directrix):

The task is to write an equation of a line given the two points: (5,45) and (10,30) That is what I got from reading the problem.

OpenStudy (itrymath):

hold on

Directrix (directrix):

What did you get for the slope of the line?

OpenStudy (itrymath):

isnt it y2-y1 over x2- x1 ??@Directrix

OpenStudy (itrymath):

@Directrix

Directrix (directrix):

Yes. Look at attached graphic.

OpenStudy (itrymath):

-3 slope

Directrix (directrix):

Yes

OpenStudy (itrymath):

so i got y=-3x+15

Directrix (directrix):

I did not get that. I'm wondering if I used the right points.

OpenStudy (itrymath):

you did

OpenStudy (itrymath):

im not right but i dont know why

Directrix (directrix):

If you show your work, I can find the error, maybe.

OpenStudy (itrymath):

lol you wont show yours <3

OpenStudy (itrymath):

okay ill show it hold on

OpenStudy (itrymath):

y-45=-3(x-5)

OpenStudy (itrymath):

then

OpenStudy (itrymath):

y=-3x+15

OpenStudy (itrymath):

its D

OpenStudy (itrymath):

i found my mistake i think.... pretty sure its D

Directrix (directrix):

>>(y - 30) = -3(x-10) y - 30 = -3x + 30 y = -3x + 30 + 30 then ....

Directrix (directrix):

You skipped a step.

OpenStudy (itrymath):

im looking over it

OpenStudy (itrymath):

so D.....

OpenStudy (itrymath):

OH my god

OpenStudy (itrymath):

B

OpenStudy (itrymath):

its B

Directrix (directrix):

Did you find the error in this: y-45=-3(x-5) Show what you got when you multiplied -3 times x and then times -5

OpenStudy (itrymath):

15

OpenStudy (itrymath):

-3x!!!

Directrix (directrix):

y-45=-3(x-5) y - 45 = -3x + 15 and then...

OpenStudy (itrymath):

+45 on both sides

OpenStudy (itrymath):

i get it ........

Directrix (directrix):

Yes

Directrix (directrix):

And, then ....

Directrix (directrix):

"Fast is fine but accuracy is everything." ~Wyatt Earp

OpenStudy (itrymath):

oh well i got a quote for you too "Invest in yourself, and never stop" ¬Warren Buffet

OpenStudy (itrymath):

and then its f(x)=-3x+60

OpenStudy (mathstudent55):

Here is another quote "Working smart saves time over working hard" - mathstudent55 Here is another way of finding the equation once you have the slope. You found out early on that the slope is -3. The equation must be of the form: y = -3x + b All you need to find is b, the y-intercept. Choose one of the given points and plug it in for x and y, and solve for b. Let's use (5, 45) y = -3x + b 45 = -3(5) + b 45 = -15 + b b = 60 Once you know b = 60, and you have the slope from before the answer is: y = -3x + 60 or in function notation: f(x) = -3x + 60

OpenStudy (itrymath):

i know.... thank you thought :)